# Adaptive Canonicalization With Application to Invariant Anisotropic Geometric Networks

### v1

@@ G.1 Sampling-based VS Optimization-based Implementation
To evaluate the trade-offs between sampling-based and our sample-and-refine (optimization-based) implementation, we conduct experiments on the TUDataset graph classification benchmarks. Tab. LABEL:tab:sampling_or_optimization reports the classification performance of TUDaset under sampling-based and optimization-based adaptive caninocalization. We see that the optimization-based implementation consistently performs better than the sampling-based one. While increasing the sampling candidates (from $1\times$ to $5\times$ or $10\times$) improves the performance, the sample-and-refine strategy is more memory-efficient than massive sampling approaches. Rather than storing and evaluating hundreds of rotation matrices simultaneously, it processes a smaller working set through iterative refinement, reducing memory pressure [Li et al., 2022]. In terms of computation time, the sampling-based method grows linearly with the number of candidates, while the optimization-based method add a small overhead to the inner steps. We see that in practice, a modest refinement (a few steps) surpasses the accuracy of large sampling budgets at a lower time, offering a better accuracy-time trade-off.
@@
G.2 Node-Level Representation Anisotropic Nonlinear Filters
We introduce the adaptive canonicalization applied to spectral graph neural networks for learning graph-level representation in Sec. 3.1 and App. E.1. Importantly, the adaptive canonicalization can also be applied to node-level representation. We describe this construction below.
In this setting, we continue to consider a classification task with $D$ classes. For each class $d$, we denote its corresponding class prior $h_{d}$. The class-$d$ orientation for node representation is obtained by optimizing
$\mathbf{U}^{\square(d)}=\bigoplus_{k}{\mathbf{U}}_{k}^{\square(d)}=\underset{\mathbf{U}^{(d)}_{1},\ldots,\mathbf{U}^{(d)}_{B}}{\arg\max}\;h_{d}\left(\Psi_{d}\left(\sum_{i}\left({\mathbf{U}}^{(d)*}\phi\left(\widetilde{\mathbf{X}}^{(d)}_{1},\ldots,\widetilde{\mathbf{X}}^{(d)}_{B}\right)\right)_{k}{\mathbf{V}}_{k}^{\top}\right)\right).$
Here, the node-level representation proceeds by mapping the input signal to the spectral domain [Mallat, 2002] in a band-wise manner with an oriented basis within each band’s eigenspace, and performing a synthesis step that transforms the learned coefficients back to the node domain.
As discussed in App. D, on a square grid, each $x$ Fourier mode has a corresponding $y$ Fourier mode of the same response. Therefore, standard spectral methods are inherently isotropic as they cannot distinguish between horizontal and vertical directional information. On the other hand, adaptive canonicalization with synthesis in the node-level setting is anisotropic. Within each degenerate eigenspace, our method can learn distinct orientations, and the subsequent synthesis step maps these anisotropic spectral coefficients back to nodes. The resulting spatial operator can therefore implement any directional filter that a convolutional neural network can achieve [Shannon, 2006, LeCun and Bengio, 1998, Freeman et al., 1991, Dagès et al., 2024]. We leave a detailed exploration of such node-level tasks for future work.
However, in graph-level tasks, the canonicalized node-level embeddings can serve as the intermediate representation from which graph-level features are derived. Specifically, after applying adaptive canonicalization and synthesis at the node level, the resulting node embeddings can be aggregated through standard pooling operations to have a graph-level representation. We evaluate the node-to-graph construction on TUDataset for graph classification tasks. The results are summarized in Tab. LABEL:tab:graph_classification_by_node_representation. We see that the node-to-graph construction achieves performance closely aligned with, and in some cases approaching, that of the direct graph-level canonicalization. We attribute the slightly worse performance to the potential pooling loss.
@@ G.3 Out-of-Sample Rotation Generalization in Point Clouds
We adopt the $z/\mathcal{SO}(3)$ protocol [Esteves et al., 2018, Deng et al., 2021]: training with on-the-fly azimuthal rotations ($z$-axis) augmentation, and evaluation applies under arbitrary rotations to each test shape. In this setting, we assess out-of-sample rotation generalization by constraining training data rotations while testing on the full rotation group. The classification performance on ModelNet40 under $z/\mathcal{SO}(3)$ protocol is reported in Tab. 7. Standard PointNet and DGCNN drop sharply under this shift. Equivariant vector-neuron variants recover much of the loss, and canonicalization baselines are comparable. Our adaptive canonicalization outperforms both equivariant architecture and canonicalization baselines in both backbones.

### v2

@@ G.1 Sampling-based VS Optimization-based Implementation
To evaluate the trade-offs between sampling-based and our sample-and-refine (optimization-based) implementation, we conduct experiments on the TUDataset graph classification benchmarks. Tab. LABEL:tab:sampling_or_optimization reports the classification performance of TUDaset under sampling-based and optimization-based adaptive caninocalization. We see that the optimization-based implementation consistently performs better than the sampling-based one. While increasing the sampling candidates (from $1\times$ to $5\times$ or $10\times$) improves the performance, the sample-and-refine strategy is more memory-efficient than massive sampling approaches. Rather than storing and evaluating hundreds of rotation matrices simultaneously, it processes a smaller working set through iterative refinement, reducing memory pressure [li2022rago]. In terms of computation time, the sampling-based method grows linearly with the number of candidates, while the optimization-based method add a small overhead to the inner steps. We see that in practice, a modest refinement (a few steps) surpasses the accuracy of large sampling budgets at a lower time, offering a better accuracy-time trade-off.
@@
G.2 Node-Level Anisotropic Nonlinear Filters
We introduce the adaptive canonicalization applied to spectral graph neural networks for learning graph-level representation in Sec. 4.1 and App. E.3. The adaptive canonicalization can also be applied to node-level representation, where the node-level representation proceeds by mapping the input signal to the spectral domain [mallat2002theory] in a band-wise manner with an oriented basis within each band’s eigenspace, and performing a synthesis step that transforms the learned coefficients back to the node domain.
On a square grid, each $x$ Fourier mode has a corresponding $y$ Fourier mode of the same response. Therefore, standard spectral methods are inherently isotropic as they cannot distinguish between horizontal and vertical directional information. On the other hand, adaptive canonicalization is anisotropic and our method can learn distinct orientations. The resulting spatial operator can therefore implement any directional filter that a convolutional neural network can achieve [shannon2006communication, lecun1998convolutional, freeman1991design, dages2024metric].
In graph-level tasks, the canonicalized node-level embeddings can serve as the intermediate representation from which graph-level features are derived. Specifically, the resulting node embeddings can be aggregated through standard pooling operations to have a graph-level representation. We evaluate the node-to-graph construction on TUDataset for graph classification tasks. The results are summarized in Tab. LABEL:tab:graph_classification_by_node_representation. We see that the node-to-graph construction achieves performance closely aligned with, and in some cases approaching, that of the direct graph-level canonicalization. We attribute the slightly worse performance to the potential pooling loss.
@@ G.3 Out-of-Sample Rotation Generalization in Point Clouds
We adopt the $z/\mathcal{SO}(3)$ protocol [esteves2018learning, deng2021vector]: training with on-the-fly azimuthal rotations ($z$-axis) augmentation, and evaluation applies under arbitrary rotations to each test shape. In this setting, we assess out-of-sample rotation generalization by constraining training data rotations while testing on the full rotation group. The classification performance on ModelNet40 under $z/\mathcal{SO}(3)$ protocol is reported in Tab. 7. Standard PointNet and DGCNN drop sharply under this shift. Equivariant vector-neuron variants recover much of the loss, and canonicalization baselines are comparable. Our adaptive canonicalization outperforms both equivariant architecture and canonicalization baselines in both backbones.
@@
G.4 Ablation Studies
We conduct ablation studies on the spectral band partitioning and the choice of GSO for A-NLSF, as well as on the impact of different point cloud backbones in our anisotropic point cloud networks.
G.4.1 Spectral Band Partition
In our experiment, we adopt a dyadic partitioning scheme (see App. E.3). In this section, we conduct an ablation using a uniform partitioning of the eigenvalues with the same number of bands and report the graph classification performance in Tab. LABEL:tab:graph_classification_different_spectral_band. We see that using the dyadic partitions performs better than using the uniform partition. tion provided by dyadic bands, which could more effectively isolate band-wise unitary actions that commute with the chosen GSO. We also note that spectral band design can be realized in more flexible and expressive ways, for example, through attention as in SpecFormer [bo2023specformer]. Investigating such learned or adaptive band-selection strategies is an important direction for future work and may further strengthen our adaptive canonicalization framework.
Table 8: Graph classification performance using uniform and dyadic spectral band partitioning.
MUTAG PTC ENZYMES PROTEINS NCI1
[…]
G.4.2 Graph Shift Operator
We evaluate the graph Laplacian as an alternative GSO. Tab. LABEL:tab:graph_classification_different_gso reports the graph classification performance of A-NLSF when instantiated with the graph Laplacian versus the normalized graph Laplacian. We observe that using the normalized graph Laplacian in our method yields better performance than the graph Laplacian. We attribute this to the properties of the normalized Laplacian: (i) the normalized Laplacian removes degree-related scaling effects, leading to a comparable spectral domain across graphs with different degree distributions, and (ii) its eigenvalues lie in the fixed interval $[0,2]$, providing a controlled and interpretable frequency range, and making dyadic partitioning better aligned across different graphs.
Table 9: Graph classification performance of A-NLSF with different GSO.
MUTAG PTC ENZYMES PROTEINS NCI1
[…]
G.4.3 Point cloud backbones
Table 10: Classification results on ModelNet40 for different point cloud backbones. Results of competing methods marked with * are taken from deng2021vector, luo2022equivariant.
Accuracy
PointNet 74.7
[…]
PointNet++ 85.0
[…]
RSCNN 82.6
[…]
DGCNN 88.6
[…]
In order to assess the impact of the backbone choice on the performance of our anisotropic point cloud networks, we extended our experiments to include two additional and widely used point cloud backbones, PointNet++ [qi2017pointnet++] and RSCNN [liu2019relation], in addition to PointNet and DGCNN reported in Tab. 3. We denote the corresponding variants by AC-PointNet++ and AC-RSCNN. The ablation results are reported in Tab. 10. We see that the choice of backbone does influence the overall point cloud classification performance. However, we observe that our adaptive canonicalization framework consistently improves the classification performance across these backbones. Moreover, when comparing methods built on the same backbone (e.g., PointNet or DGCNN), our approach outperforms equivariant models, data augmentation, and standard canonicalization (see Tab. 3). This indicates that our method is robust across different point cloud backbones and can further benefit from stronger backbones when they are available.
G.5 Sensitivity Analysis
To examine the effect of different hyperparameters, we conduct a hyperparameter sensitivity study covering grid size, sinusoidal period, noise level, and hidden dimension. For each hyperparameter, we swept over a range of values while keeping all other settings fixed. The results of the sensitivity analysis are summarized in Fig. 2. Overall, we observe that our method is reasonably robust. For grid size and sinusoidal period, performance remains stable across the tested ranges. For the noise level, small to moderate noise leads to similar performance, with a degradation only when the noise becomes large enough that it effectively corrupts the underlying structure of the data. For the hidden dimension, small dimensions impact the performance, but performance stabilizes once we enter a standard regime of model capacity.
Figure 2: Hyperparameter sensitivity with respect to grid size, noise level, and hidden dimension.
G.6 Truncation Canonicalization with a pretrained classifier
We introduce in App. E.5 an application of our adaptive canonicalization on truncation prior maximization. We now illustrate the applicability of this setup with a pretrained image classifier. Specifically, we take a ResNet-18 [he2016deep] pretrained on ImageNet [deng2009imagenet]. We freeze the backbone, and train only the classifier using the CIFAR-10 or CIFAR-100 [krizhevsky2009learning] training set. The experiment is conducted with ten independent runs, and the resulting image classification performance is reported in the Tab. LABEL:tab:truncation_image. We see that truncation-based prior maximization improves classification performance over the standard vanilla baseline. This implies that our method enables the model to adaptively select a canonical truncation that enhances downstream performance. In addition, we observe that the selected canonical crops tend to tightly focus on the main object while discarding background and irrelevant context. It matches the intuition behind our prior maximization: by optimizing over the truncation family, the model chooses a representative transformed image that best aligns with its prior over the class. This experiment demonstrates that our adaptive canonicalization framework can be instantiated with a truncation symmetry and benefit from off-the-shelf pretrained models. It also highlights the potential of transformation families as a practical way to improve pretrained models via adaptive canonicalization.
Table 11: Image classification accuracy on CIFAR-10 and CIFAR-100 using a ResNet18 pretrained on ImageNet, with and without truncation canonicalization.
CIFAR-10 CIFAR-100
[…]
G.7 Computational requirement comparison
Tab. LABEL:tab:compute_budget the training time per epoch with the number of parameters. We see that A-NLSF uses a similar number of parameters as the other methods and fewer than the spectral method. Its computational requirements are comparable to the other methods and does not rely on a significantly larger training budget than the competing methods.
Table 12: Running time per epoch(s)/number of parameters.
MUTAG PTC ENZYMES PROTEINS NCI1
MLP 0.07/105K 0.10/114K 0.13/125K 0.37/129K 1.01/134K
GCN 0.40/116K 0.66/120K 0.81/137K 1.92/142K 5.84/149K
GAT 0.62/138K 0.87/149K 0.96/154K 2.34/159K 4.93/167K
GIN 0.14/105K 0.37/106K 0.52/107K 0.94/106K 1.97/121K
ChebNet 0.79/185K 1.25/189K 1.72/191K 3.64/217K 11.52/245K
FA+GIN 0.57/120K 1.04/123K 1.35/126K 2.55/130K 4.31/142K
OAP+GIN 0.22/105K 0.39/104K 0.57/109K 1.21/110K 2.36/124K
A-NLSF 0.44/132K 0.89/140K 1.29/145K 2.20/148K 4.37/151K
G.8 Training stability
To quantify the training stability of our method, we track the canonical rotations of a subset of 1500 randomly chosen training examples in the point cloud classification experiment. At each epoch, we measure the mean geodesic distance on $\mathcal{SO}(3)$ between the canonicalizations between consecutive epochs. Fig. 3 reports the mean geodesic distance between epochs. We observe that this distance decreases during the training and then remains stable, indicating that the canonical representatives stabilize with no rapid switching.
Figure 3: Mean geodesic distance on $\mathcal{SO}(3)$ between the canonicalizations between consecutive epochs.
G.9 Canonicalized point clouds
Fig. 4 shows the canonicalized point clouds for the chair class in the point cloud classification experiment. We randomly select 20 examples from this class and visualize them after applying the optimal transformations. We observe that the examples in this class share a similar orientation after canonicalization.
Figure 4: The canonicalized point clouds for the chair class.
G.10 ShapeNet Part Segmentation
Table 13: Part segmentation performance on the ShapeNet part dataset. The metric is reported with the average category mean IoU Results of competing methods marked with * are taken from deng2021vector, kaba2023equivariance.
PointNet 62.3
DGCNN 78.6
VN-PointNet 72.8
VN-DGCNN 81.4
[…]
To expand our experimental study on point cloud data, we further conduct experiments on the ShapeNet part segmentation benchmark [yi2016scalable]. The dataset consists of 16 shape categories annotated with a total of 50 parts, where each category is labeled with between two and six parts. Note that our prior maximization adaptive canonicalization method was naturally suited to classification tasks. Extending it to the segmentation task is not trivial, as the segmentation task requires predicting a label for each point in the point cloud, and applying prior maximization to each point would be computationally inefficient. Therefore, in order to adapt our adaptive canonicalization to the segmentation task, we consider the adaptive canonicalization with the minimal entropy prior summing over nodes. This modification preserves the spirit of the adaptive canonicalization while making it compatible with per-point prediction. Tab. 13 reports the segmentation performance. For the PointNet backbone, we see that, similar to the point cloud classification task, the entropy-based adaptive canonicalization yields advantageous segmentation performance compared to equivariant architectures and standard canonicalization baselines. For the DGCNN backbone, our method attains performance comparable to equivariant architectures while outperforming existing canonicalization methods. These results demonstrate that our approach has potential beyond classification. Note that one of the main contributions of our work is to construct continuous and symmetry-respecting models. In the entropy prior adaptive canonicalization, the continuity property is not straightforward. We plan to investigate the continuity properties of this adapted approach in future work.